# Inferring Resource-Oriented Intentions using LLMs for Static Resource Leak Detection

### v3

@@ Abstract
Resource leaks, caused by resources not being released after acquisition, often lead to performance issues and system crashes. Existing static detection techniques rely on mechanical matching of predefined resource acquisition/release APIs and null-checking conditions to find unreleased resources, suffering from both (1) false negatives caused by the incompleteness of predefined resource acquisition/release APIs and (2) false positives caused by the incompleteness of resource reachability validation identification.
To overcome these challenges, we propose InferROI, a novel approach that leverages the exceptional code comprehension capability of large language models (LLMs) to directly infer resource-oriented intentions (acquisition, release, and reachability validation) in code. InferROI first prompts the LLM to infer involved intentions for a given code snippet, and then incorporates a two-stage static analysis approach to check control-flow paths for resource leak detection based on the inferred intentions. We evaluate the effectiveness of InferROI in both resource-oriented intention inference and resource leak detection. Experimental results on the DroidLeaks and JLeaks datasets demonstrate InferROI achieves promising bug detection rate (59.3% and 64.8%) and false alarm rate (18.6% and 24.0%). Compared to three industrial static detectors, InferROI detects 14~45 and 167~503 more bugs in DroidLeaks and JLeaks, respectively. When applied to real-world open-source projects, InferROI identifies 26 unknown resource leak bugs, with 7 of them being confirmed by developers. Finally, manual annotation indicated that InferROI achieved a precision of 74.6% and a recall of 81.8% in intention inference, covering more than 60% resource types involved in the datasets. The results of an ablation study underscores the importance of combining LLM-based inference with static analysis.
@@ I Introduction
In this paper, we propose InferROI, a novel approach that leverages large language models (LLMs) to boost static resource leak detection. Our intuition is that LLMs have been trained on massive code and documentation corpus and thus have exceptional code comprehension capability [9, 10, 11, 12], suggesting their potential for inferring resource-oriented intentions (i.e., the functionalities of resource acquisition/release and resource reachability validation) in any given code snippet. InferROI consists of two stages, i.e., resource-oriented intention inference and lightweight resource leak detection. (1) For a given code snippet, InferROI first prompts the LLM to directly infer the involved resource acquisition, release, and reachability validation intentions, without relying on any prior knowledge of resource API pairs. The intentions generated by LLMs (which are described in natural language) are then formalized into formal expressions, so as to facilitate the subsequent static analysis stage. (2) InferROI then incorporates a two-stage static analysis approach to check control-flow paths for resource leaks based on the inferred intentions. In particular, InferROI first identifies potential leak-risky paths based on the inferred resource acquisition/release APIs and then prunes leak-risky paths with unreachable resources based on the inferred reachability validation, so as to reduce false alarms of resource leak detection.
While our approach is general and not limited to specific programming languages, we currently implement InferROI for Java given its prevalence. We conduct an extensive evaluation to assess the efficacy of InferROI in detecting resource leaks. Firstly, we apply InferROI to 86 and 784 bugs sourced from the DroidLeaks [8] and JLeaks [13] datasets, respectively, to evaluate its effectiveness in resource leak detection. The results reveal InferROI demonstrates a bug detection rate of 59.3% with a false alarm rate of 18.6% for DroidLeaks, and a bug detection rate of 64.8% with a false alarm rate of 24.0% for JLeaks. Compared to three industrial static detectors, InferROI detects 14~45 and 167~503 more bugs in DroidLeaks and JLeaks, respectively, while maintaining a comparable false alarm rate. Secondly, we employ InferROI for resource detection in real-world open-source projects. Our experiments on 100 methods uncover 26 previously unknown resource leaks, 7 of which are confirmed by developers as of the submission time. Finally, manual annotation for resource-oriented intentions indicated that InferROI achieved a precision of 74.6% and a recall of 81.8% in intention inference, covering 19 (67.9%) and 221 (60.1%) resource types involved in DroidLeaks and JLeaks, respectively. The results of an ablation study underscores the importance of combining LLM-based inference with static analysis.
@@ II Motivating Example
2. Identifying resource reachability validation more comprehensively. The validation of resource reachability is a crucial factor in determining the occurrence of resource leaks. In the fixed version depicted in Figure 1, the if-condition at line 185 serves to validate the reachability of the acquired HTTP client resource before its release at line 186. If this if-condition for reachability validation is disregarded, a false alarm will be triggered in the false branch of the if-statement, where the acquired resource is actually unreachable.
@@ III Approach
Figure 2 represents the approach overview of InferROI. For a given code snippet, InferROI first instantiates the prompt template to instruct the LLM in inferring the involved resource acquisition, release, and reachability validation intentions, which does no require any prior information on resource API pairs. The intentions generated by LLMs (which are described in natural language) are then formalized into formal expressions, so as to facilitate the subsequent static analysis stage. By aggregating these inferred intentions, InferROI then proceeds to apply a two-stage static analysis algorithm to analyze the control-flow paths extracted from the code. The algorithm first identifies potential leak-risky paths based on the inferred resource acquisition/release APIs, and then prunes leak-risky paths with unreachable resources based on the inferred reachability validation across these paths on sibling branches, so as to reduce false alarms of resource leak detection.
@@ III-A Resource-Oriented Intentions
We focus on the following three resource-oriented intentions involved in code, which play a crucial role in resource management.
Resource Acquisition. Statements acquire resources (e.g., locks or database connections), which are normally API invocations and involved with objects/variables of resources (e.g., lock or db). To formally express the intentions of resource acquisition, we introduce the notation $\bm{ACQ(var,lineno)}$, where $var$ denotes the variable storing the acquired resource, and $lineno$ indicates the line number containing the statement responsible for acquiring the resource.
Resource Release. Likewise, statements release previously acquired resources. We formalize the intentions of resource release as $\bm{REL(var,lineno)}$, where $var$ corresponds to the variable storing the resource to be released, and $lineno$ indicates the line number of the statement responsible for performing the release operation.
Resource Reachability Validation. Conditional statements aim to validate whether acquired resources remain reachable within specific branches. To formalize the intentions of resource reachability validation, we employ the notation $\bm{VAL(var,lineno)}$, where $var$ represents the variable storing the resource being validated, and $lineno$ indicates the line number of the statement responsible for performing the validation operation.
@@ III-B LLM-Based Intention Inference
To infer resource-oriented intentions present in code, we first prompt the large language model (LLM) with designed templates and then formalize the LLM-generated answers into formal expressions for the subsequent analysis.
@@ III-B1 Feeding LLM with Prompt
For a given code, we construct a prompt using the template depicted in Figure 3. This prompt is then fed into the LLM (i.e., GPT-4 used in this work) for intention inference. The prompt template is designed following the best practices for prompt engineering by OpenAI [22], consisting of three main components: Task Description & Instructions, Output Format Specification, and Code Placeholder.
@@
Task Description & Instructions component introduces the background (the first sentence) and outlines the instructions (the remaining sentences) to be executed by the LLM. The prompt includes five sequential instructions, with the first two serving as leading instructions to guide the LLM in smoothing the reasoning chain from a basic starting point. In the leading instructions, the LLM is prompted to perform type inference for the involved objects and determine which types represent leakable resources. Based on the analysis, the LLM is then required to infer the three types of resource-oriented intentions by the follow-up three instructions. Throughout the analysis and inference process, the LLM leverages its extensive background knowledge to understand the code contexts and execute the instructions effectively.
Output Format Specification aims to normalize the format of identified intentions, facilitating subsequent parsing of the outputs.
@@
III-B2 Extracting Formal Expressions
To extract formal expressions of the intentions identified by the LLM, we employ matching patterns that correspond to the output format specification specified in the template. Given an LLM-generated answer, we utilize regular expressions such as “line <line number>: <API call> acquires <resource variable> resource” to extract the line number, API call, and resource variable involved in resource acquisition intentions. Similar regular expressions can be used to extract information related to resource release intentions and reachability validation intentions. We then use the extracted information to instantiate the formal expressions of the three types of resource-oriented intentions.
Example 3.3. From the answer illustrated in Figure 4, the three highlighted lines are matched by the matching patterns. Subsequently, the three formal intention expressions are extracted and instantiated as $ACQ({\texttt{\footnotesize client}},167)$, $REL({\texttt{\footnotesize client}},186)$, and $VAL({\texttt{\footnotesize client}},185)$.
@@ III-C Static Resource Leak Detection
We then propose a lightweight static analysis based on the inferred resource-oriented intentions to detect resource leaks in code.
@@ III-C1 Enumerating Control-Flow Paths
Similar to most existing resource leak detection methods [3, 4, 6, 7], our detection approach is also built upon control-flow analysis. We construct control-flow graphs (CFGs) for the code to detect and extract pruned control-flow paths from these graphs.
Given a method-level code snippet, we begin by parsing it into an abstract syntax tree (AST) and then proceed to construct a CFG by traversing the AST nodes and inserting various control-flow edges between the code statements. The procedure is implemented based on the top of an established CFG construction tool [24]. Within the CFG, certain nodes, such as if-statements and switch-statements, give rise to multiple subsequent control-flow branches. These particular nodes are referred to as branch-nodes within the CFG. Next, we enumerate the paths within the CFG, considering the paths between the entry-node and exit-node. To ensure the efficiency of the enumeration process, we employ two pruning operations.
@@
Example 3.4. For the buggy version in Figure 1, there is only one control-flow path that sequentially executes from line 160 to line 185. For the fixed version in Figure 1, there are two control-flow paths branched by the if-statement of reachability validation at line 185. The two control-flow paths represented in line numbers are: [160-185, 186, 187-190] and [160-185, 187-190].
@@ III-C2 Detecting Resource Leaks
Stage 1: Single-Path Analysis. The algorithm conducts an iterative process, sequentially processing each $path$ in $\mathcal{P}$ and traversing the nodes within $path$ (lines 2-17). During this node traversal, a descriptor counter $rd\_counter$ keeps track of the $ACQ$ and $REL$ intentions involving the resource $res$ within $path$ (lines 3-11). If $rd\_counter$ is greater than 0, it indicates that $path$ is a potential leak-risky path for $res$, and as a result, the risky status of $path$ (i.e., $path.risky$) is initialized as $true$ (lines 12-16).
@@
Intuitively, addressing this problem involves specifically analyzing each branch-node intended for reachability validation and identifying its unreachable branch. However, the task is complicated by the diverse implementations of validation APIs, as discussed in Section II. Some APIs (e.g., isClosed()) return true to indicate that the resources are reachable, while others (e.g., isActive()) indicate the opposite. To overcome this challenge, our approach utilizes a novel heuristic that collectively analyzes all branches of a branch-node to assess whether it introduces false alarms. If affirmative, such branch-nodes are designated as false-alarm-introducing (FAI) nodes. These FAI nodes are characterized by its one branch being completely leak-risky, while the others are completely not leak-risky. A branch is considered completely leak-risky when all the paths belonging to it exhibit a risky status of $true$. The completely leak-risky branch is the unreachable branch, and the paths within it are FP leak-risky paths for $res$. The insight behind this decision is that if the other branch (the completely not leak-risky branch) is the unreachable one, it suggests the presence of more severe functional bugs rather than resource leaks, such as null pointer errors.
@@ IV Implementation
The current implementation of InferROI utilizes the widely adopted GPT-4 as the LLM and sets the model temperature as 0 to remove generation randomness. During the static resource leak detection, we take into account certain Java-specific language features, such as try-with-resources statements. For the try-with-resources statement, the acquired resources are automatically closed, and there is no need for manual resource closure [25]. To address this, we incorporate a simple rule-based post-processing step to filter out any false alarms introduced by the try-with-resources statements.
@@ V Evaluation
We conduct a comprehensive set of experiments to evaluate the effectiveness of InferROI in detecting resource leaks by employing the DroidLeaks dataset [8] and the JLeaks dataset [13]. InferROI typically takes an average of 3 to 8 seconds to process each code snippet.
First, we apply InferROI to detect resource leaks in the buggy-version and fixed-version code snippets in the datasets, comparing its bug detection rate and false alarm rate against those of baseline detectors (RQ1). By doing so, we assess whether InferROI can surpass the performance of the baseline detectors in accurately identifying resource leaks. In the next part of our evaluation (RQ2), we explore the efficacy of InferROI in detecting resource leaks when applied to real-world open-source projects. This investigation enables us to gauge InferROI’s performance under real-world conditions. Finally, we manually annotate the resource-oriented intentions presented in code to measure the extent to which InferROI successfully infers these intentions (RQ3.a), and conduct an ablation study to analyze and compare the contributions of two key components of InferROI (RQ3.b).
@@
  • RQ1 (Effectiveness in Leak Detection): How effectively can InferROI detect resource leaks in the DroidLeaks and JLeaks datasets? Can it outperform baseline detectors in this regard?
  • RQ2 (Effectiveness in Project Scanning): How effective is InferROI in detecting previously-unknown resource leak bugs when applied to open-source projects?
  • RQ3.a (Accuracy in Intention Inference): To what degree can InferROI accurately infer the resource-oriented intentions in code?
  • RQ3.b (Ablation Study): What is the contribution of the intention inference and static analysis in InferROI?

### v4

# Boosting Static Resource Leak Detection via LLM-based Resource-Oriented Intention Inference

@@ Abstract
Resource leaks, caused by resources not being released after acquisition, often lead to performance issues and system crashes. Existing static detection techniques rely on mechanical matching of predefined resource acquisition/release APIs and null-checking conditions to find unreleased resources, suffering from both (1) false negatives caused by the incompleteness of predefined resource acquisition/release APIs and (2) false positives caused by the incompleteness of resource reachability validation identification. To overcome these challenges, we propose InferROI, a novel approach that leverages the exceptional code comprehension capability of large language models (LLMs) to directly infer resource-oriented intentions (acquisition, release, and reachability validation) in code. InferROI first prompts the LLM to infer involved intentions for a given code snippet, and then incorporates a two-stage static analysis approach to check control-flow paths for resource leak detection based on the inferred intentions.
We evaluate the effectiveness of InferROI in both resource-oriented intention inference and resource leak detection. Experimental results on the DroidLeaks and JLeaks datasets demonstrate InferROI achieves promising bug detection rate (59.3% and 62.5%) and false alarm rate (18.6% and 19.5%). Compared to three industrial static detectors, InferROI detects 14~45 and 149~485 more bugs in DroidLeaks and JLeaks, respectively. When applied to real-world open-source projects, InferROI identifies 29 unknown resource leak bugs (verified by authors), with 7 of them being confirmed by developers. In addition, the results of an ablation study underscores the importance of combining LLM-based inference with static analysis. Finally, manual annotation indicated that InferROI achieved a precision of 74.6% and a recall of 81.8% in intention inference, covering more than 60% resource types involved in the datasets.
@@ I Introduction
In this paper, we propose InferROI, a novel approach that leverages large language models (LLMs) to boost static resource leak detection. Our intuition is that LLMs have been trained on massive code and documentation corpus and thus have exceptional code comprehension capability [9, 10, 11, 12], suggesting their potential for inferring resource-oriented intentions (i.e., the code statements for resource acquisition/release and resource reachability validation) in any given code snippet. InferROI consists of two stages, i.e., resource-oriented intention inference and lightweight resource leak detection. (1) For a given code snippet, InferROI first prompts the LLM to directly infer the involved resource acquisition, release, and reachability validation intentions, without relying on any prior knowledge of resource API pairs. The intentions generated by LLMs (which are described in natural language) are then converted into structured expressions, so as to facilitate the subsequent static analysis stage. (2) InferROI then incorporates a two-stage static analysis approach to check control-flow paths for resource leaks based on the inferred intentions. In particular, InferROI first identifies potential leak-risky paths based on the inferred resource acquisition/release APIs and then prunes leak-risky paths with unreachable resources based on the inferred reachability validation, so as to reduce false alarms.
While our approach is general and not limited to specific programming languages, we currently implement InferROI for Java given its prevalence. We conduct an extensive evaluation to assess the efficacy of InferROI in detecting resource leaks. Firstly, we apply InferROI to 86 and 784 bugs sourced from the DroidLeaks [8] and JLeaks [13] datasets, respectively, to evaluate its effectiveness in resource leak detection. The results reveal InferROI demonstrates a bug detection rate of 59.3% with a false alarm rate of 18.6% for DroidLeaks, and a bug detection rate of 62.5% with a false alarm rate of 19.5% for JLeaks. Compared to three industrial static detectors, InferROI detects 14~45 and 149~485 more bugs in DroidLeaks and JLeaks, respectively, while maintaining a comparable false alarm rate. Secondly, we employ InferROI for resource detection in real-world open-source projects. InferROI uncovers 29 previously unknown resource leaks from 200 methods, 7 of which are confirmed by developers as of the submission time. In addition, an ablation study underscores the importance of combining LLM-based inference with static analysis. The generalizability of InferROI with different LLMs is confirmed by an evaluation involving Llmam3-8B and Gemma2-9B. Finally, manual annotation indicates that InferROI achieves a precision of 74.6% and a recall of 81.8% in intention inference, covering 19 (67.9%) and 221 (60.1%) resource types involved in DroidLeaks and JLeaks, respectively.
@@ II Motivating Example
2. Identifying resource reachability validation more comprehensively. The validation of resource reachability is a crucial factor in determining the occurrence of resource leaks. Note that the meanings between reachable/reachability and the related terms (e.g., open/closed/valid) are different. Open/closed/valid describe the status of a resource itself; while reachable/reachability indicate whether the open status of a resource can be propagated to a certain path, branch, or statement. For example, in the fixed version in Figure 1, the if-condition at line 185 validates the reachability of the resource “HTTP client” by checking whether the object of client is null. As a result, the client resource is actually unreachable on the false branch of this if-condition, that is, the open status of the client resource cannot reach the false branch. If a resource leak detection technique ignores the reachability of the client resource, it would cause a false alarm in the false branch of this if-condition. Therefore, reachability is essential for reducing false alarms in static resource leak detection.
@@ III Approach
Figure 2 represents the approach overview of InferROI. For a given code snippet, InferROI first instantiates the prompt template to instruct the LLM in inferring the involved resource acquisition, release, and reachability validation intentions, which does no require any prior information on resource API pairs. The intentions generated by LLMs (which are described in natural language) are then converted into structured expressions, so as to facilitate the subsequent static analysis stage. By aggregating these inferred intentions, InferROI then proceeds to apply a two-stage static analysis algorithm to analyze the control-flow paths extracted from the code. The algorithm first identifies potential leak-risky paths based on the inferred resource acquisition/release APIs, and then prunes leak-risky paths with unreachable resources based on the inferred reachability validation across these paths on sibling branches, so as to reduce false alarms of resource leak detection.
@@ III-A Resource-Oriented Intentions
We focus on three resource-oriented intentions involved in code, which play a crucial role in resource management.
Resource Acquisition. Statements acquire resources (e.g., locks or database connections), which are normally API invocations and involved with objects/variables of resources (e.g., lock or db). To structurally express the intentions of resource acquisition, we introduce the notation $\bm{ACQ(var,lineno)}$, where $var$ denotes the variable storing the acquired resource, and $lineno$ indicates the line number containing the statement responsible for acquiring the resource.
Resource Release. Likewise, statements release previously acquired resources. We structurally represent the intentions of resource release as $\bm{REL(var,lineno)}$, where $var$ corresponds to the variable storing the resource to be released, and $lineno$ indicates the line number of the statement responsible for performing the release operation.
Resource Reachability Validation. Conditional statements aim to validate whether acquired resources remain reachable within specific branches. To structurally represent the intentions of resource reachability validation, we employ the notation $\bm{VAL(var,lineno)}$, where $var$ represents the variable storing the resource being validated, and $lineno$ indicates the line number of the statement responsible for performing the validation operation.
@@ III-B LLM-Based Intention Inference
To infer resource-oriented intentions present in code, we first prompt the large language model (LLM) with designed templates and then convert the LLM-generated answers into structured expressions for the subsequent analysis.
@@ III-B1 Feeding LLM with Prompt
For a given code, we construct a prompt using the template depicted in Figure 3. This prompt is then fed into the LLM (e.g., GPT-4) for intention inference. The prompt template is designed following the best practices for prompt engineering by OpenAI [22], consisting of three main components: Task Description & Instructions, Output Format Specification, and Code Placeholder.
@@
The Task Description & Instructions component consists of several sequential instructions designed based on the chain-of-thought (CoT) idea [23]. It begins with three leading instructions to introduce the overall task goal and guide the LLM in forming a smooth reasoning chain from a basic starting point, specifically by performing type inference for the involved objects and determining which types represent leakable resources. Based on the analysis, the LLM is then required to infer the three types of resource-oriented intentions by the follow-up three instructions. Throughout the analysis and inference process, the LLM leverages its extensive background knowledge to understand the code contexts and execute the instructions effectively.
Output Format Specification aims to specify the format of identified intentions, facilitating subsequent parsing.
@@
III-B2 Extracting Structured Expressions
To extract structured expressions of the intentions identified by the LLM, we employ matching patterns that correspond to the output format specification specified in the template. Given an LLM-generated answer, we utilize regular expressions such as “line <line number>: <API call> acquires <resource variable> resource” to automatically extract the line number, API call, and resource variable involved in resource acquisition intentions. Similar regular expressions can be used to extract information related to resource release intentions and reachability validation intentions. We then use the extracted information to instantiate the structured expressions of the three types of intentions.
Example 3.3. From the answer illustrated in Figure 4, the three highlighted lines are matched by the matching patterns. Subsequently, the three structured intention expressions are extracted and instantiated as $ACQ({\texttt{\footnotesize client}},167)$, $REL({\texttt{\footnotesize client}},186)$, and $VAL({\texttt{\footnotesize client}},185)$.
@@ III-C Static Resource Leak Detection
We then propose a lightweight static analysis based on the inferred resource-oriented intentions to detect resource leaks.
@@ III-C1 Enumerating Control-Flow Paths
Similar to most existing resource leak detection methods [3, 4, 6, 7], our detection approach is also built upon control-flow analysis. We construct control-flow graphs (CFGs) for the code to detect and extract pruned control-flow paths from these graphs. Within the CFG, nodes that generate multiple subsequent control-flow branches, such as if-statements and switch-statements, are referred to as branch-nodes. Next, we enumerate the paths within the CFG, considering the paths between the entry-node and exit-node. To ensure the efficiency of the enumeration process, we employ two pruning operations.
@@
Example 3.4. For the buggy version in Figure 1, there is only one control-flow path that sequentially executes from line 160 to line 185. In the fixed version shown in Figure 1, the if-statement for reachability validation at line 185 introduces two control-flow paths, corresponding to the line ranges [160-185, 186, 187-190] and [160-185, 187-190].
@@ III-C2 Detecting Resource Leaks
Stage 1: Single-Path Analysis. The algorithm conducts an iterative process, sequentially processing each $path$ in $\mathcal{P}$ and traversing the nodes within $path$ (lines 2-17). During the traversal, a descriptor counter $rd\_counter$ keeps track of the $ACQ$ and $REL$ intentions involving $res$ within $path$ (lines 3-11). If $rd\_counter$ is greater than 0, it indicates that $path$ is a potential leak-risky path for $res$, resulting in the risky status of $path$ (i.e., $path.risky$) is initialized as $true$ (lines 12-16). When no released API is identified, there is no $REL$ item in the inferred intention expression. For example, when there is an $ACQ$ in a path but no corresponding $REL$, the rd_counter will be 1, marking the path as risky (Line 12).
@@
Intuitively, addressing this problem involves specifically analyzing each branch-node intended for reachability validation and identifying its unreachable branch. However, the task is complicated by the diverse implementations of validation APIs, as discussed in Section II. Some APIs (e.g., isClosed()) return true to indicate that the resources are reachable, while others (e.g., isActive()) indicate the opposite. To overcome this challenge, our approach utilizes a novel heuristic that collectively analyzes all branches of a branch-node for reachability validation to assess whether it introduces false alarms. If affirmative, such branch-nodes are designated as false-alarm-introducing (FAI) nodes. These FAI nodes are characterized by its one branch being completely leak-risky, while the others are completely not leak-risky. A branch is considered completely leak-risky when all the paths belonging to it exhibit a risky status of $true$. The completely leak-risky branch is the unreachable branch, and the paths within it are FP leak-risky paths for $res$. The insight behind this decision is that if the other branch (the completely not leak-risky branch) is unreachable, it suggests the presence of more severe functional bugs rather than resource leaks, such as null pointer errors.
To illustrate, consider the if-condition node (line 185) in the Fixed Version in Figure 1. It validates the client resource and then releases it. When one branch is completely leak-risky (marked in Stage 1 due to not containing client.close()) and the other is completely not leak-risky (marked in Stage 1 due to containing client.close()), we can essentially determine that this if-condition is a FAI node and its completely leak-risky branch is the unreachable branch. Otherwise, if client.close() is called in an unreachable branch (e.g., if(client==null){client.close();}), it would result in a more severe and obvious null-pointer exception.
@@ IV Implementation
The current implementation of InferROI utilizes the widely adopted GPT-4 as the LLM and sets the model temperature as 0 to remove generation randomness. For RQ4, we use Meta-Llama-3-8B-Instruct [25] and gemma-2-9b-it [26] versions available on the Hugging Face Platform, and run the model inference with the Transformers library [27]. To align with the temperature setting (i.e., temperature=0) of GPT-4, we use greedy search decoding for both models, ensuring there is no randomness. The experiments are run on a workstation with 4 AMD EPYC 9554 64-Core CPUs, 512 GB of memory, and a single Nvidia H100 (80GB) GPU. For the CFG construction in static resource leak detection, given a method-level code snippet, we begin by parsing it into an abstract syntax tree (AST) and then proceed to construct a CFG by traversing the AST nodes and inserting various control-flow edges between the code statements. The procedure is implemented based on the top of an established CFG construction tool [28]. Since some resources are intentionally kept open and returned for further use, we adopt a rule to exclude the leak reports where the inferred resource objects returned in return-statements, thereby reducing false positives. In addition, we take into account certain Java-specific language features, such as try-with-resources statements. For the try-with-resources statements, the acquired resources are automatically closed, and there is no need for manual resource closure [29]. To address this, we incorporate a rule-based post-processing step to filter out false alarms caused by try-with-resources statements. Specifically, for each reported resource leak, we traverse the AST of the examined method to determine if the leaked resource is acquired within a try-with-resources statement. If so, the reported leak is omitted.
@@ V Evaluation
We conduct comprehensive experiments to evaluate the effectiveness of InferROI in detecting resource leaks by employing the DroidLeaks dataset [8] and the JLeaks dataset [13].
First, we apply InferROI to detect resource leaks in the code snippets from DroidLeaks and JLeaks datasets, comparing its effectiveness against baseline detectors (RQ1). Next, we evaluate the effectiveness of InferROI in detecting previously unknown resource leaks in real-world open-source projects (RQ2). We then conduct an ablation study to evaluate the contributions of intention inference and static analysis in InferROI (RQ3). Subsequently, we investigate the generalizability of integrating InferROI with different LLMs and compare their effectiveness (RQ4). Finally, we manually annotate the resource-oriented intentions in code to measure how effectively InferROI infers these intentions (RQ5).
@@
  • RQ1 (Effectiveness in Leak Detection): How effectively can InferROI detect resource leaks in existing benchmarks?
  • RQ2 (Effectiveness in Project Scanning): How effective is InferROI in detecting previously unknown resource leak bugs when applied to open-source projects?
  • RQ3 (Ablation Study): What are the contributions of intention inference and static analysis in InferROI?
  • RQ4 (Generalizability with Different LLMs): How generalizable is InferROI when integrated with different LLMs?
  • RQ5 (Accuracy in Intention Inference): How accurately can InferROI infer resource-oriented intentions in code?